OpenStudy (anonymous):

x^2-20x=64= (factorize)

OpenStudy (anonymous):

hmm.. pretty sure you made a mistake in those "=" signs

OpenStudy (anonymous):

subtract 64 both side so that the equation can equal to 0

OpenStudy (anonymous):

x^2-20x+64=?

OpenStudy (anonymous):

x^2-20x-64=0

OpenStudy (anonymous):

Ah.. there ya go. It's a "+" instead of "=" so the factor is (x-16) (x-4)

OpenStudy (anonymous):

(x-4)(x-16)

OpenStudy (anonymous):

you prob have a typo

OpenStudy (anonymous):

(x-10)^2-160=0

OpenStudy (anonymous):

is this a diff question?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no